### v1

# ICP-4D: Bridging Iterative Closest Point and LiDAR Panoptic Segmentation

Abstract
Dominant paradigms for 4D LiDAR panoptic segmentation are usually required to train deep neural networks with large superimposed point clouds or design dedicated modules for instance association. However, these approaches perform redundant point processing and consequently become computationally expensive, yet still overlook the rich geometric priors inherently provided by raw point clouds. To this end, we introduce ICP-4D, a simple yet effective training-free framework that unifies spatial and temporal reasoning through geometric relations among instance-level point sets. Specifically, we apply the Iterative Closest Point (ICP) algorithm to directly associate temporally consistent instances by aligning the source and target point sets through the estimated transformation. To stabilize association under noisy instance predictions, we introduce a Sinkhorn-based soft matching. This exploits the underlying instance distribution to obtain accurate point-wise correspondences, resulting in robust geometric alignment. Furthermore, our carefully designed pipeline, which considers three instance types—static, dynamic, and missing—offers computational efficiency and occlusion-aware matching. Our extensive experiments across both SemanticKITTI and panoptic nuScenes demonstrate that our method consistently outperforms state-of-the-art approaches, even without additional training or extra point cloud inputs.
**footnotetext: Corresponding author.
1 Introduction
Autonomous driving systems heavily depend on precise perception of dynamic 3D environments with geometrically accurate measurements captured by LiDAR sensors. In particular, LiDAR segmentation plays a crucial role in enabling comprehensive point-wise understanding. In 3D environments, LiDAR panoptic segmentation [19, 32] originates from semantic and instance segmentation, which jointly identify class categories and instance IDs. Along with the recent significant advances in frame-wise perception, 4D LiDAR panoptic segmentation [2] has emerged to capture temporal relationships in a holistic manner. In this context, ensuring temporal continuity is a principal purpose in 4D LiDAR panoptic segmentation, even under instance motions.
With its potential benefits, as shown in Figure 1, the mainstream of prior works can be categorized into three discrete paradigms: 1) IoU-based association, 2) Query-propagated association, and 3) Detect & Track association. One solution [2, 20, 63, 54, 16], superimposing point clouds, is widely adopted for establishing 4D spatio-temporal volume. It enables consistent instance tracking over time by measuring the overlapped regions (see Figure 1(a)). However, despite its simplicity, stacking point clouds (i.e., increasing the number of points per batch) leads to a quadratic growth in computational overhead, which is an inherent challenge in point cloud processing [28, 22, 50, 51].
An alternative direction [30, 1, 29] seeks to mitigate this issue by treating each scan independently, as illustrated in Figure 1(b) and 1(c). They are primarily derived from the two major paradigms of multi-object tracking, tracking-by-detection [7, 49, 58, 36] and query propagation [55, 59, 31], to assign consistent IDs for instances. Both approaches differ substantially in training strategies, yet commonly suffer from error accumulation [59, 41] and reliance on costly and dataset-dependent training. Consequently, we argue that this suboptimal design fails to fully exploit the strong perception capability of 3D panoptic networks, which naturally raises the following question: Can instance association be achieved efficiently using only a 3D panoptic network?
Figure 1: Comparison of 4D LiDAR panoptic segmentation methods: (a) IoU-based, (b) Query-propagated, (c) Detect & Track, and (d) ICP-4D. Instead of relying on training with large-scale point clouds, ours achieves reliable association in a fully training-free manner. Refer to caption indicates methods that require training and Refer to caption denotes frozen network.
To answer this, we propose Iterative Closest Point (ICP)-based strategy for reliable instance matching, even without any training process. ICP is widely used for point cloud registration [46, 57, 11], but its applicability beyond that has been largely underexplored. Here, we first integrate ICP into the 4D LiDAR panoptic segmentation framework. This allows explicit mapping through estimated transformation between instance segment pairs in successive scans. While this mapping works well with complete segment pairs, noisy predictions hinder finding corresponding points via nearest-neighbor approach, as it ignores the global instance structure. To handle this, we propose Sinkhorn-based soft matching that strengthens the robustness against outliers with instance-aware correspondences. Specifically, we represent each instance’s point set as a probability distribution. Then, the correspondence problem is cast as a Sinkhorn Optimal Transport (OT) problem [12], which computes an entropy-regularized transport plan between the two distributions. This plan identifies point pairs that minimize the transportation cost, thereby establishing soft correspondences.
ICP demonstrates competitive performance in instance association. However, its iterative mechanism and the physical limitations of LiDAR sensors introduce remaining challenges, a marginal increase in runtime and vulnerability to occlusion. Given these limitations, we classify instances into three categories: static, dynamic, and missing. For associating static instances, we employ its point set statistics (i.e., mean and covariance) considering the rigid property. This significantly reduces the number of matching candidates and preserves computational efficiency. Moreover, to compensate for occluded instances, we present a memory bank to store prediction results within a short temporal window.
To wrap it up, we named our overall pipeline as ICP-4D and demonstrate its effectiveness through extensive experiments on the SemanticKITTI [4] and nuScenes [13] benchmarks. Note that ICP-4D achieves superior performance over existing state-of-the-art methods and ranks 1st on the SemanticKITTI test leaderboard, even under the minimal setting of a single scan and without any additional training.
The contributions of this work are summarized as follows:
  • We introduce ICP-4D that is a novel training-free 4D LiDAR panoptic segmentation pipeline, flexibly integrated with a wide range of 3D panoptic segmentation networks.
  • We propose a Sinkhorn-based soft matching strategy to enhance the robustness of point correspondences, and we delicately design instance state-conditioned solutions to effectively mitigate time complexity and occlusion.
  • Extensive experiments in challenging environments clearly demonstrate the superiority of ICP-4D, supported by in-depth analysis.
2 Related Work
2.1 LiDAR Panoptic Segmentation
LiDAR panoptic segmentation simultaneously handles two different basis tasks, semantic and instance segmentation. In early, proposal-based methods [27, 34, 52] filter the background points with Mask R-CNN [14]’s semantic head, then group the thing points into instances. In contrast, proposal-free methods [15, 37, 62, 25] independently predict semantic prediction and class-agnostic instance segments via end-to-end training manners. 4D LiDAR panoptic segmentation extends these frameworks to sequential scans, requiring instance IDs to remain consistent over time.
In terms of association strategy, predominant works [60, 2, 20, 63, 54, 16] aggregate consecutive point clouds (e.g., 2-4 scans) and group instances in overlapping regions to ensure temporal continuity. Alternatively, CA-Net [29] associates instances via feature-wise similarity, followed by the outputs of 3D panoptic network. Mask4D [30] preserves identity consistency across scans by iteratively updating the track queries in an end-to-end manner. Despite these attempts, we observe significant issues regarding the computational cost and limited adaptability (see Sec. 1). Instead, we propose ICP-4D that is a flexible, efficient, and well-balanced instance association through direct geometric registration.
2.2 Multi-Object Tracking (MOT)
MOT is an indispensable task for autonomous driving systems, enabling the tracking of all detected objects across scans. Tracking-by-detection (TBD) paradigm [7, 49, 58, 36] has become the de facto standard due to its simplicity and the rapid progress of detectors. Here, the motion model (e.g., Kalman Filter [18]) plays a key role by estimating future object states to maintain tracking continuity. The emergence of DETR [64] has recently drawn increasing attention to the end-to-end paradigm [55, 59, 31], which performs detection and tracking in parallel using query-based tracking. These paradigms are not confined to 2D images but are increasingly embraced in 3D autonomous driving scenarios [48, 61, 35, 56] as well. However, despite sharing objectives with MOT, 4D LiDAR panoptic segmentation differs fundamentally, as it requires holistic point-level scene understanding instead of bounding boxes. Although ICP-4D follows the TBD paradigm, it operates directly on point-level panoptic outputs, enabling instance association grounded in the geometric structure of the underlying point clouds.
2.3 Iterative Closest Point (ICP)
ICP [6] is one of the most well-established and widely adopted algorithm for point-set registration [9, 24, 39]. Over the years, many variants of the vanilla ICP have been proposed to address its limitations, and have been widely adopted in both traditional engineering [8, 38, 40, 53] and deep learning–based applications [46, 57, 11]. Nowadays, ICP-Flow [26] proposes a compelling scene flow prediction framework that inspired this work. They only consider point-wise motion estimation between two temporally adjacent scans by leveraging well-separated instance clusters. In this paper, we introduce ICP into the context of the 4D LiDAR panoptic segmentation, ensuring the temporal continuity of instances across the entire scans. Remarkably, our Sinkhorn-based soft matching enables instance-centric association, even when predictions are noisy, imprecise, or fragmented.
3 Method
ICP-4D primarily focuses on ensuring temporal continuity of the instances in an unsupervised manner. To handle this, we adopt Iterative Closest Point (ICP) [6] to associate instances through their geometric consistency over time. In this section, we first provide the introductory ICP (Sec. 3.1), then go over the technical details of our methods (Sec. 3.2).
3.1 Preliminary
Iterative Closest Point (ICP). Conventionally, for point cloud registration, ICP estimates the optimal transformation by iteratively minimizing the difference between source and target point sets. Given two point sets $\mathcal{P}_{src}\in\mathbb{R}^{n\times 3}$ and $\mathcal{P}_{dst}\in\mathbb{R}^{m\times 3}$, it starts with the initial guess of rigid transformation ($\mathbf{R}^{(0)},\mathrm{t}^{(0)}$), where the rotation $\mathbf{R}\in\text{SO}(3)$ and translation $\mathrm{t}\in\mathbb{R}^{3}$, which satisfies the following formula:
$\mathcal{P}_{dst}\approx\mathbf{R}\mathcal{P}_{src}+\mathrm{t}\;.$ (1)
Here, in order to compute the alignment error, pairwise correspondences are established using a nearest-neighbor search algorithm (e.g., KD-tree [5]). Classical ICP adopts $l_{2}$ distance as the error metric between the matched point pairs. Ultimately, the optimal rigid transformation is obtained by iteratively solving a series of least-squares problems as:
[…]
to update the rotation $\mathbf{R}$ and translation $\mathrm{t}$ until convergence. With this theoretical background, we investigate how ICP contributes to consistent instance ID assignment across frames while enhancing robustness against outliers.
3.2 ICP-4D
Our proposed ICP-4D is developed on top of a pre-trained 3D LiDAR panoptic segmentation model [15, 54], which is specialized in frame-wise perception. Through this model, we obtain the per-point instance and semantic prediction $(\mathcal{I}_{i},\mathcal{S}_{i})$ for a given point cloud with $N$ points, where $\mathcal{S}_{i}\in\{1,...,C\}$, $\mathcal{I}_{i}\in\mathbb{N}_{0}$, and $C$ is the number of classes. Prior to the association stage, we construct a set of matching candidates between $t$ and t1t-1 times by pairing source and destination instances as follows:
$\mathcal{G}_{\text{src}}\times\mathcal{G}_{\text{dst}}=\{(a_{i},b_{j})\;|\;a_{i}\in\mathcal{G}_{\text{src}},\;b_{j}\in\mathcal{G}_{\text{dst}}\},$ (3)
where $\mathcal{G}_{\text{src}}$ and $\mathcal{G}_{\text{dst}}$ are respectively denoted as the sets of unique instance IDs in the source and destination. Then, we derive the filtered set $\mathcal{M}=\mathcal{G}^{\prime}_{\text{src}}\times\mathcal{G}^{\prime}_{\text{dst}}$ by keeping only the pairs in $\mathcal{G}_{\text{src}}\times\mathcal{G}_{\text{dst}}$ that share the same semantic class. In this study, we categorize instances into dynamic, static, and missing types to facilitate effective instance association across frames.
3.2.1 Static Instance
Under the assumption that both the center and the shape of a non-moving rigid object remain invariant, we filter static instances based on their centers and further refine them using covariance cues to improve selection fidelity. Given the matching candidates $\mathcal{M}$, we first represent each instance by the set of 3D points belonging to it. For all instances in $\mathcal{G}^{\prime}_{\text{src}}$ and $\mathcal{G}^{\prime}_{\text{dst}}$, we denote their corresponding point sets as $P(\mathcal{G}^{\prime}_{\text{src}})$ and $P(\mathcal{G}^{\prime}_{\text{dst}})$, respectively. Then, we compute the center of each point set:
$\mu_{s}={1\over|P(\mathcal{G}^{\prime}_{\text{src}})|}\sum_{u\in P(\mathcal{G}^{\prime}_{\text{src}})}\!\!\!u,\quad\mu_{d}={1\over|P(\mathcal{G}^{\prime}_{\text{dst}})|}\sum_{v\in P(\mathcal{G}^{\prime}_{\text{dst}})}\!\!\!v,$ (4)
where $u,v\in\mathbb{R}^{3}$ and indicate the world coordinates of points. We keep only the pairs satisfying $\mu_{d}-\mu_{s}||_{2}<\tau_{center}$, where $\tau_{center}$ is the distance threshold for identifying non-moving objects based on center consistency. Although this naive approach performs well under ideal conditions, it suffers when applied to noisy predicted point sets, resulting in unsuitable pair associations. Thus, we introduce an additional covariance-based constraint that measures the similarity of the spatial spread and geometric structure of the instance point sets. Specifically, the covariance of each point set is computed as follows:
${\sigma}_{s}=\mathrm{Cov}(P(\mathcal{G}^{\prime}_{\text{src}})),\quad{\sigma}_{d}=\mathrm{Cov}(P(\mathcal{G}^{\prime}_{\text{dst}})).$ (5)
To assess shape consistency, we measure the similarity between the two covariance matrices and use the predefined threshold $\tau_{cov}$ to control the discrepancy between source and destination point set:
[…]
where $\text{tr}(\cdot)$ and $\|\cdot\|_{F}$ denote the matrix trace and the Frobenius norm. Then, based on the obtained similarity scores, we choose the highest score for each source instance to find the best-matching pairs while avoiding duplicate associations.
Figure 2: Illustration of Sinkhorn-based soft matching. (a) Nearest neighbor-based matching aligns each source point with its closest counterpart, focusing only on local point proximity. (b) In contrast, correspondences are computed by transport plan $\mathcal{Q}$, enabling instance-aware matching that respects the global geometry of each instance point set.
3.2.2 Dynamic Instance
After filtering static instances, the remaining candidates $\mathcal{G}_{\text{src}}^{\prime\prime}$ and $\mathcal{G}_{\text{dst}}^{\prime\prime}$ are fed into the ICP loop that is the core mechanism of our framework that enables instance association without any supervised training. In the vanilla ICP loop, the first step is finding point pairs to compute the distance by utilizing the nearest-neighbor algorithm [5]. While ICP has a powerful ability of registering point cloud sets, determining the corresponding points under noisy instance predictions potentially causing erroneous matches. Therefore, we propose a Sinkhorn-based soft matching scheme that facilitates stable correspondence estimation within the ICP framework.
Sinkhorn-based Soft Matching. Sinkhorn-based soft matching starts from regarding the point cloud as a probability distribution rather than discrete set of individual points. Given two point sets $P(\mathcal{G}_{\text{src}}^{\prime\prime})=\{u_{i}\}^{I-1}_{0}$ and $P(\mathcal{G}_{\text{dst}}^{\prime\prime})=\{v_{j}\}^{J-1}_{0}$, where $I$ and $J$ denote the number of points in each segment, we formulate correspondence estimation as an Optimal Transport (OT) problem [12]. Specifically, we seek a transport plan $\mathcal{Q}$ that minimizes the cost of transforming one distribution into the other as follows:
$\text{OT}(P(\mathcal{G}^{\prime\prime}_{\text{src}}),\;P(\mathcal{G}^{\prime\prime}_{\text{dst}})):=\underset{\mathcal{Q}}{\operatorname*{min}}\langle\mathcal{Z},\;\mathcal{Q}\rangle_{F}-\epsilon H(\mathcal{Q}),$ (7)
[…]
Returning to the ICP procedure, we follow the standard process, initializing transformations with a histogram-based method [26] to reduce sensitivity to initial estimates and subsequently running iterative refinement to get the final rigid transformation. After obtaining the transformation, we determine whether two instances correspond by evaluating the IoU between their transformed point sets. Specifically, points with errors smaller than the predefined threshold $\tau_{dist}$ are treated as inliers. Then, in order to filter valid matching candidates, we evaluate alignment accuracy between source and destination instances by IoU scores as follows:
$\mathcal{M}^{dyn}=\begin{cases}\text{matched},&\!\!\!\!\text{if}~\langle\mathcal{F}(P(\mathcal{G}^{\prime\prime}_{\text{src}})),P(\mathcal{G}^{\prime\prime}_{\text{dst}})\rangle\geq\tau_{\text{iou}},\\ \text{rejected},&\!\!\!\!\text{otherwise},\end{cases}$ (8)
where $\langle\cdot\rangle$ denotes the IoU calculation, and $\mathcal{F}$ applies the estimated transformation obtained by ICP.
Option: Bijective assignment. A naive approach is to pair each source instance with its most similar target instance; however, it fails to enforce one-to-one correspondence. Thus, we derive a transformation-aware matching cost that integrates the ICP results—rotation $\mathbf{R}$, translation $\mathrm{t}$, and the IoU score $s_{\text{iou}}$. Formally, the matching cost $\mathcal{C}$ of dynamic matching pairs $(i,j)\in\mathcal{M}^{dyn}$ is defined as:
$\mathcal{C}_{ij}$ $=\gamma_{t}c_{ij}^{(t)}+\gamma_{r}\,c_{ij}^{(r)}+\gamma_{s}\,c_{ij}^{(s)},$ (9)
$c_{ij}^{(t)}=\|\mathrm{t}_{ij}\|_{2},$ $c_{ij}^{(r)}=\|\theta_{ij}\|_{2},\quad c_{ij}^{(s)}=\|1-s_{iou}\|_{2},$
where $\theta$ denotes the axis–angle rotation vector obtained from $\mathbf{R}$. We normalize each cost term to comparable scales before weighting so that $\gamma_{t}$, $\gamma_{r}$, and $\gamma_{s}$ control their relative contributions. By minimizing $\mathcal{C}$ with the Hungarian algorithm [21], we obtain a bijective assignment that considers the geometric alignment.
3.2.3 Missing Instance
Occlusions stemming from the inherent limitations of LiDAR sensor often break temporal continuity beyond two consecutive scans. To mitigate this, we introduce a memory bank $\mathcal{B}$ that caches the world coordinates, semantic labels, and instance IDs of unmatched objects. Owing to bounded memory capacity, $\mathcal{B}$ keeps only entries observed within a short window of the previous $w_{\text{mem}}$ scans. If $\mathcal{B}$ is nonempty, we retrieve previously stored instance information and associate using the dynamic instance matching process.
4 Experiments
Table 1: Quantitative results on SemanticKITTI benchmark [4]. We highlight the best scores in bold and the underline indicates the second-best scores. Note that * represents that the results are obtained from our implementation following the official training setup.
Assoc. Training Validation set Test set
Method # Scan LSTQ $_{\text{assoc}} S$$_{\text{cls}}$ IoUSt IoUTh LSTQ $_{\text{assoc}} S$$_{\text{cls}}$ IoUSt IoUTh
(a) IoU-based Association
4D-PLS [2] 4 62.7 65.1 60.5 65.4 61.3 56.9 56.4 57.4 66.9 51.6
4D-StOP [20] 4 67.0 74.4 60.3 65.3 60.9 63.9 69.5 58.8 67.7 53.8
Eq-4D-StOP [63] 4 70.1 77.6 63.4 66.4 67.1 67.0 72.0 62.4 69.1 60.9
Mask4Former [54] 2 70.5 74.3 66.9 67.1 66.6 68.4 67.3 69.6 72.7 65.3
black!10black       (b) Query-propagated Association
Mask4D [30] 1 71.4 75.4 67.5 65.8 69.9 64.3 66.4 62.2 69.9 52.2
black!10black       (c) Detect & Track Association
KPConv [43]+PP [23]+MOT [47] 1 46.3 37.6 57.0 64.2 54.1 38.0 25.9 55.9 66.9 47.7
KPConv+PP+SFP [33] 1 46.0 37.1 57.0 64.2 54.1 38.5 26.6 55.9 66.9 47.7
black!10black DS-Net[15] 1 - - - 63.5 64.6 62.0 - - 60.6 66.9 52.0
[…]
Mask4Former 1 - - - 67.5 65.2 70.6 - - 70.5 72.7 67.4
[…]
Table 2: Quantitative results on panoptic nuScenes [13].
Method # Scan LSTQ $_{\text{assoc}} S$$_{\text{cls}}$ IoUSt IoUTh
(a) IoU-based Association
4D-StOP [20] 4 60.5 62.5 58.6 74.4 54.9
Eq-4D-StOP [63] 4 67.3 73.7 61.5 76.4 58.7
Mask4Former [54] 2 67.6 63.0 72.5 76.6 66.3
black!10black       (b) Query-propagated Association
Mask4D [30] 1 58.5 54.1 63.1 75.5 55.7
black!10black       (c) Detect & Track Association
Mask4Former 1 - - 71.0 76.0 65.9
[…]
4.1 Experiment Setup
Datasets. We choose two representative benchmarks. SemanticKITTI [4] captures point clouds using a 64-beam LiDAR at 10 Hz and provides point-wise annotations for 19 semantic classes (8 Things / 11 Stuff). In contrast, panoptic nuScenes [13] uses a 32-beam LiDAR operating at 2 Hz and offers point-wise annotations for 16 semantic classes (10 Things / 6 Stuff). This contrast serves as a valuable testbed for assessing generalization across dense (SemanticKITTI) and sparse (nuScenes) LiDAR settings.
Evaluation Metrics. To evaluate both tracking and segmentation quality, we adopt the widely used LSTQ metric [2], defined as $\text{LSTQ}=\sqrt{S_{cls}\times S_{assoc}}$. Here, $S_{cls}$ measures how accurately semantic labels are assigned and $S_{assoc}$ evaluates how consistently instance identities are maintained throughout the sequence, independently of semantic accuracy. However, $S_{assoc}$ is computed only from true positive associations weighted by IoU scores. Thus, when prediction masks are stable and yield high IoU scores, ID switch causes a sharper drop in this score. Consequently, while ICP-4D primarily influences $S_{assoc}$, the LSTQ must be taken into account as a meaningful indicator of temporal consistency.
Implementation Details. Our proposed ICP-4D can be incorporated into any 3D LiDAR panoptic segmentation network. For fair comparison, we basically follow two baseline’s configuration, DS-Net [15] and Mask4Former [54]. Since Mask4Former is originally developed for 4D panoptic segmentation, we train it using the official settings with the input modified to a single scan instead of multiple sweeps. For instance association, thresholds for filtering instance pairs are set to $\tau_{center}=\tau_{cov}=\tau_{dist}=0.1$ and $\tau_{iou}=0.2$. Additionally, the short window $w_{\text{mem}}$ is set to $3$ for missing instances, the regularization parameter for optimal transport is set to $\epsilon=0.2$, and the number of ICP iterations is fixed to 30. All experiments are conducted on a single RTX A6000.
Figure 3: Qualitative comparison on SemanticKITTI validation set. We visualize the association results over five consecutive scans for both the baselines and our ICP-4D. Different colors represent different instances. The dotted boxes zoom in for clear comparison. For clarity, we use circled numbers (①–⑤) to represent the frame indices.
Figure 4: Comparison of computational efficiency. We illustrate the efficiency trade-offs between memory usage (left) & runtime (right) and performance across different methods. Yellow triangles show LSTQ scores for each model. Red dashed line denotes our memory & runtime, and yellow dashed line marks our LSTQ score.
4.2 Quantitative Analysis
4.2.1 4D LiDAR Panoptic Segmentation Results
SemanticKITTI. Table 4 reports the experimental results on SemanticKITTI validation and hidden test set. Here, we clearly observe two insights. First, ICP-4D achieves the highest LSTQ score across all data split regardless of the number of input scans. Specifically, while (a) IoU-based association approaches rely on 22-44 more input point clouds, our ICP-4D achieves superior performance using only a single scan, showing improvements of up to 13.2 percent points in $S_{assoc}$ and 10.0 percent points in LSTQ. Furthermore, by comparing with the learning-based approaches (b) and (c), another noticeable aspect is that ICP-4D consistently outperforms recent state-of-the-art methods on both the validation and test sets, whereas Mask4D and CA-Net drop in performance on the test set. This highlights the robustness of our direct geometric registration with raw point clouds, rather than relying on training-dependent feature association.
nuScenes. Table 4 shows the 4D panoptic results on panoptic nuScenes dataset. Consistent with the observations on the SemanticKITTI dataset, ICP-4D delivers notably stronger instance association performance, outperforming existing methods in terms of LSTQ. It is important to note that baselines generally suffer when the LiDAR scans are captured with a long interval. In such scenarios, multi-scan aggregation benefits from denser temporal support and thus exhibit strong performance. For example, EQ-4D-StOP with 4 aggregated scans surpasses Mask4Former with 2 scans by a substantial margin in the association metric. Nevertheless, our ICP-4D demonstrates that even without accessing multiple scans, a single-scan input is sufficient to achieve competitive performance. This result implies that ICP-4D remains highly effective even under sparse LiDAR configurations.
4.2.2 Comparison of Efficiency
Next, we examine computational efficiency in terms of memory usage and runtime. For each method, we report the peak GPU memory consumption and the average runtime measured over the entire validation set. Figure 4 illustrates the results, supporting that ICP-4D offers a most well-balanced trade-off between performance and computational overhead. First, when compared with the IoU-based association approaches, we observe that ICP-4D requires 60.7%–79.4% less memory and achieves a 26.1%–45.6% runtime efficiency improvement, all while delivering superior LSTQ performance. While Mask4D exhibits high efficiency in runtime, it inevitably incurs substantial memory overhead due to its transformer-based architecture design. Lastly, CA-Net is computationally efficient; however, it fails to obtain reliable association performance. This gap is already observable on the validation set and becomes even more pronounced on the hidden test set, as highlighted in blue in Table 4.
4.3 Qualitative Analysis
In this section, we deeply analyze the effectiveness of ICP-4D qualitatively by presenting the visualization results in Figure 3. In the first row, there are two bicyclists riding in a line. Both ground truth and our ICP-4D consistently distinguish the two bicyclists across time. However, Mask4Former struggles to preserve identity consistency when objects are spatially adjacent or overlapping due to its IoU-based association. Furthermore, Mask4D employs learnable tracking queries for association, which can lead to confused tracking queries when objects share similar visual or spatial features. Then, in the second row, we emphasize the effectiveness of our memory bank in handling missing instances. Even when the 3D panoptic model fails to detect the person colored in green at the second time step, the memory bank successfully retrieves the missing instance and preserves its identity across scans. In contrast, Mask4Former experiences ID switching when it fails to detect instance due to the absence of the overlapped region. Consequently, these observations underscore the importance of explicit geometric matching in our approach and a well-structured association pipeline.
Table 3: Ablation on module design.
Method Metric
[…]
1 - - - 70.1 73.4
2 - - 71.3 75.4
3 - - 71.4 75.6
4 - 72.0 76.8
5 72.7 (+2.6) 78.3 (+4.9)
Table 4: Ablation study on different types of instances.
Instance Type Metric
[…]
- - 35.4 18.5 1.06
- - 71.8 76.7 3.98
- 72.0 76.8 2.06
72.7 78.3 2.35
-41.0%-48.2%
4.4 Ablation Study
Module Design. Table 3 shows the effects of our proposed methods. Exp. #1 directly applies the vanilla Iterative Closest Point (ICP) for associating instances, serving as a baseline to examine the feasibility of ICP in an unsupervised manner. Despite its strong matching capability, ICP is highly sensitive to the quality of initial alignment. In Exp. #2, we observe a little performance gain when applying histogram-based initialization proposed by [26], which stabilizes the initial alignment. Notably, our Sinkhorn-based soft matching brings further improvements without the influence of initial pose (see Exp. #3). The key takeaway is that instance-aware correspondence to alleviate the instability of local point matching is highly beneficial for obtaining reliable matching points between noisy instance segments, and as a result helps to estimate an accurate transformation. Exp. #4 demonstrates that coupling these components yields a complementary effect, leading to more robust instance association and improved overall performance. Lastly, Exp. #5 demonstrates the effectiveness of incorporating a memory bank for handling occluded and missing instances, resulting in superior performance.
Effect of Different Types of Instances. To validate our carefully designed pipeline with respect to different instance types, Table 4 presents the association results and average inference time per a single scan. First, we employ the association strategy designed for static instances (Sec. 3.2.1) to matching candidates over all detected objects. Despite requiring minimal computation, only a small subset of instances remain consistently tracked, 18.5 percent points indicating that static cues alone are insufficient for reliable association. In contrast, relying solely on dynamic instance association (Sec. 3.2.2) achieves high performance, but is computationally expensive. Thus, combining these strategies reduces the runtime by 48.2% while simultaneously improving performance, as it significantly narrows the matching candidates for dynamic instance association. Consequently, our full ICP-4D enhances the tracking continuity without sacrificing computational efficiency.
Table 5: Ablation on matching strategy.
Method LSTQ $_{\text{assoc}} S$$_{\text{cls}}$
Hungarian Matching 72.3 77.6 67.5
Greedy Matching 72.7 78.3 67.5
Table 6: Ablation on DBSCAN instance refinement.
Method LSTQ $_{\text{assoc}} S$$_{\text{cls}}$ Inf. Time (s)
w/o DBSCAN [10] 71.8 75.8 67.5 0.84
w/ DBSCAN 72.7 78.3 67.5 2.35
[…]
Analysis of DBSCAN Instance Refinement. In this specific experiment of integrating ICP-4D with Mask4Former, we apply DBSCAN clustering [10] as done in the original Mask4Former setting. Table 6 summarizes the results, where we derive two major insights. First, ICP-4D consistently surpasses existing state-of-the-art methods, both with and without the additional clustering step, achieving at least a 0.4 percentage point gain in LSTQ. Furthermore, we discover that the increase of runtime primarily results from the post-processing step. Without this step, ICP-4D delivers the most efficient runtime of 840 ms compared to all other methods, with only a marginal loss in performance. These observations clearly highlight that ICP-4D offers both high efficiency and strong performance, and can further benefit from improvements in the backbone network.
[…]
Effect of Sinkhorn-based Soft Matching. The Sinkhorn-based soft matching forms instance-aware point correspondence between source and target points. Owing to this strategy, it allows accurate error measurements albeit with outliers, by not being overwhelmed by the local proximity. Figure 5 visualizes the point-wise correspondences obtained from two consecutive scans, with the current scan serving as the source and the previous scan as the target. For this analysis, we extract each point segment at the initial stage of ICP, as this stage retains its geometric structure intact. As shown in the highlighted region, the nearest neighbor matching is biased toward the closest point, ignoring the underlying point-wise relationships of the instance. In contrast, applying Sinkhorn-based soft matching on the baseline maintains the geometric structure, yielding correspondences that align consistently with the instance’s motion direction.
Class-wise Performance Analysis. In this section, we further analyze the effectiveness of ICP-4D w.r.t. the class over the SemanticKITTI validation set. We categorize classes with a large number of observed points (car, truck, other-vehicle) into $S_{assoc}^{b}$, whereas classes with naturally small point counts (person, bicyclist, motorcyclist, motorcycle, bicycle) are grouped into $S_{assoc}^{s}$. As shown in Figure 6, ICP-4D excels at maintaining instance consistency not only for vehicle classes but also for challenging small object classes, outperforming the baselines. Additionally, it is worth noting that all classes except bicycle exhibit favorable association performance (see Figure 7). This is attributed to the robustness of ICP-4D under scenarios where object sizes and occurrence frequencies vary significantly.
Figure 6: Comparison of association quality by size group.
Figure 7: Class-wise association performance.
5 Conclusion
In this paper, we introduce a novel instance association strategy ICP-4D for 4D LiDAR panoptic segmentation, the first training-free framework tailored for a single scan setting. Our proposed Sinkhorn-based soft matching forming instance-aware correspondence enables ICP procedure to precisely align between two consecutive instance point segments, thereby improving temporal continuity. Additionally, we systematically design overall pipeline, underscoring the instance types to alleviate the inherent challenges of ICP and LiDAR sensors. Through experimental evaluations on both SemanticKITTI and nuScenes, our ICP-4D showcases its superiority not only on the validation sets but also on the hidden test sets, demonstrating robustness across various sensor configuration. Ultimately, our approach offers a promising direction for 4D perception by solely leveraging strong 3D panoptic models to support temporal understanding.
Limitations. Although ICP-4D shows substantial promise in advancing 4D perception, several limitations persist that warrant further investigation, particularly in real-time deployment, handling partial LiDAR observations, and dependence on camera parameters. We provide additional discussions in the supplementary materials.
Supplementary Material
Appendix A Implementation Details
A.1 Baseline Details
In this section, we provide a detailed overview of the strong baselines considered in our comparison: 4D-PLS [2] (CVPR’21), 4D-Stop (ECCVW’22) [20], EQ-4D-Stop (ICCV’23) [63], CA-Net (R-AL’22) [29], Mask4D (R-AL’23) [30], and Mask4Former (ICRA’24) [54]. Note that we report the official results as stated in the original papers, and reproduce the results when formal reports are not available. The following paragraph details the experimental settings and notable modifications for each baseline.
First, 4D-PLS, 4D-Stop, and EQ-4D-Stop follow a two-stage design, where instance tracklets are first generated independently for all scans and then stitched across consecutive frames. Accordingly, we measure the inference time of each stage separately and report their average as the overall runtime. Second, Mask4Former is originally developed for 4D LiDAR panoptic segmentation with superimposed point clouds. To establish a strong baseline for our experiments, we train it as a 3D panoptic network without any modifications, except setting the number of input scans to one. Third, CA-Net employs contrastive learning for instance association on top of an off-the-shelf 3D panoptic segmentation network. For a fair comparison, we integrate the single scan Mask4Former into the CA-Net framework. Since CA-Net requires point-wise instance features for contrastive learning, we extract them by computing the dot product between instance queries and instance masks. Lastly, official panoptic nuScenes results are unavailable for any methods except 4D-Stop and EQ-4D-Stop. Therefore, we additionally reproduce the remaining baselines using their official repositories, following the experimental protocol of EQ-4D-StOP.
A.2 Metrics
In this section, we provide detailed explanations of the metrics employed for evaluating association performance in our experiments. While LSTQ is the most widely used metric, we additionally report two complementary metrics, MOTSA and PTQ, to provide a more comprehensive assessment.
LSTQ [2] is delicately designed to evaluate 4D panoptic segmentation quality. First, to assess the semantic segmentation performance over the classes $C$, the classification score Scls is defined as:
$S_{cls}={1\over{|C|}}\sum_{c=1}^{C}{|\text{TP}_{c}|\over{|\text{TP}_{c}|+|\text{FN}_{c}|+|\text{FP}_{c}|}}={1\over{C}}\sum_{c=1}^{C}IoU(c),$ (10)
where TP, FN, and FP denote the true positives, false negatives, and false positives, respectively. Furthermore, the association score Sassoc, which is responsible for measuring temporal continuity, is defined as:
$S_{assoc}={1\over{|I|}}\sum_{t\in I}{1\over|{gt_{id}(t)}|}\sum_{s\in S,\;s\cap t\neq 0}TPA(s,t)IoU(s,t),$ (11)
where $I$ and $S$ refer to the ground-truth and predicted segments. Note that $TPA$ represents the points that maintain consistent identities across scans, regardless of class. Consequently, LSTQ is computed as the geometric mean of the classification score Scls and association score Sassoc.
MOTSA & sMOTSA [45] are mask-based variants of the widely used Multi-Object Tracking Accuracy (MOTA) metric, replacing bounding-box overlaps with mask IoU to better evaluate segmentation-level tracking performance. Formally,
$\text{MOTSA}={{|TP|-|FP|-|IDSW|}\over{|M|}},$ (12)
where $|IDSW|$ denotes the total number of ID switches, and $|M|$ is the number of ground truth segmentation masks. In addition, sMOTSA replaces the true positive (TP{TP}) with soft true positive ($sTP$), obtained by accumulating the mask IoUs over all matched predictions.
PTQ & sPTQ [17] are formulated to satisfy the fundamental requirement that panoptic tracking metrics must jointly capture segmentation accuracy and tracking consistency over time. Thus, they incorporate an ID-switch penalty into the original Panoptic Quality (PQ) measure, as follows:
$\text{PTQ}={1\over{|C|}}\sum_{c=1}^{C}{\sum_{(s,t)\in TP_{c}}IoU(s,t)-{|IDSW_{c}|}\over{|TP_{c}|+{1\over 2}|FP_{c}|+{1\over 2}|FN_{c}|}}.$ (13)
Furthermore, sPTQ leverages the soft version of IDSW (sIDSW) instead of IDSW. This soft variant aggregates the IoU scores for segments in which the predicted instance ID $ID$ changes:
$sIDSW_{c}=\{IoU(s,\;t)\;\mid\;(s,\;t)\in TP_{c}\land ID_{s}\neq ID_{t}\}.$ (14)
Algorithm 1 Instance association process of ICP-4D
0:  Instance predictions $\mathcal{I}$, semantic predictions $\mathcal{S}$, memory bank $\mathcal{B}$
1:  for time step $t\in\{1,2,\dots,T-1\}$ do
2:   Construct the candidate set $\mathcal{M}=\mathcal{G}^{\prime}_{\text{src}}\times\mathcal{G}^{\prime}_{\text{dst}}$ filtered by semantic predictions $\mathcal{S}$. // Static instances (Sec. 3.2.1)
3:   Associate instances using the center $\mu$ of each point set (Eq. 4).
4:   Refine the matched pairs using the covariance descriptor $\sigma$ (Eq. 6). // Dynamic instances (Sec. 3.2.2)
5:   Perform dynamic instance association using ICP enhanced with Sinkhorn-based soft matching. // Missing instances (Sec. 3.2.3)
6:   if $\mathcal{B}$ is not empty then
7:    Apply the dynamic matching procedure (Lines 5) to missing instances.
8:   end if
9:   Update the memory bank $\mathcal{B}$.
10:  end for
10:  Consistent instance IDs across the entire sequence.
Appendix B Method Details
B.1 Pipeline Details
The overall association pipeline of our proposed ICP-4D is summarized in Algorithm 1. We apply ICP-4D sequentially across all scans within each sequence, and reset the memory bank $\mathcal{B}$ at the beginning of every new sequence.
B.2 Complexity Analysis
In this section, we provide a computational complexity comparison between Sinkhorn-based soft matching and nearest neighbor-based matching. As described in Section.3.2.2, we utilize the entropy-regularized optimal transport problem [12] to obtain instance-aware point correspondences. Before analyzing the regularized form, we briefly review the optimal transport problem. Given two probability vectors $r\in\mathbb{R}_{+}^{n}\;\text{and}\;c\in\mathbb{R}_{+}^{m}$, the original optimal transport problem [44] aims to find the optimal transport plan $\mathcal{Q}\in U(r,\;c)$. The entries of $\mathcal{Q}$ represent the amount of mass transported from the $i$-th element of $r$ to the $j$-th element of $c$. The transport polytope $U$ is defined as:
$U(r,\;c)$ $=$ (15)
$\{Q\in$ $\mathbb{R}^{n\times m}_{+}|\sum_{j}Q_{ij}=r_{i},\;\forall i,\;\sum_{i}Q_{ij}=c_{j},\;\forall j\}.$
Let $\mathcal{Z}$ denote the transport cost matrix. The optimal transport plan is then computed as the Frobenius inner product between $\mathcal{Q}$ and $\mathcal{Z}$:
$\text{OT}(r,c):=\underset{\mathcal{Q}}{\operatorname*{min}}\langle\mathcal{Z},\;\mathcal{Q}\rangle_{F}.$ (16)
This optimization is a linear program, for which the network simplex [3] is the most widely used exact solver. In the symmetric case $n=m$, its worst-case computational complexity scales as $O(n^{3}\log n)$. However, due to its heavy computational cost, adding an entropic regularization is widely adopted to accelerate optimal transport. By incorporating the Shannon entropy $H(\mathcal{Q}):=-\sum_{i,j}\mathcal{Q}_{ij}\log\mathcal{Q}_{ij}$ into Eq. 16, we obtain the entropy-regularized optimal transport formulation. To handle marginal constraints, we introduce dual variables $f\in\mathbb{R}^{n}$ and $g\in\mathbb{R}^{m}$ and form the corresponding Lagrangian. Then, by taking the first-order optimality condition, we find that optimal solution $\mathcal{Q}$ has matrix form $Q=\text{diag}(u)K\text{diag}(v)$, where $u_{i}=e^{f_{i}/\epsilon}$, $v_{j}=e^{g_{j}/\epsilon}$, and $K_{ij}=e^{-Z_{ij}/\epsilon}$. Since $u$ and $v$ must satisfy the marginal constraints, the entropic regularization transforms the classical optimal transport problem from a linear program into a matrix scaling problem. Thus, when $n=m$, the computational complexity is reduced to $O(n^{2})$. However, the nearest neighbor algorithm operates in $O(n\log n)$ on average and up to $O(n^{2})$ in the worst case in practice, as implemented in the scikit-learn library (e.g., using KD-Tree [5]). Consequently, our Sinkhorn-based solution is theoretically comparable to the nearest neighbor algorithm, and the wall-clock time shows only a marginal difference while achieving higher performance (see Table A.1).
Table A.1: Comparison Between Nearest Neighbor and Sinkhorn-based Matching. We report the results of our full ICP-4D when integrating each correspondence strategy.
w/ Nearest Neighbor (KD-Tree) w/ Sink.-based Soft Matching
LSTQ Inf. Time (s) LSTQ Inf. Time (s)
72.1 2.01 72.7 2.35
Table A.2: SemanticKITTI [4] Test Set Leaderboard (Top 5). mAQ and mIoU denote the Sassoc and Scls, respectively.
Rank User LSTQ $\uparrow$ mAQ $\uparrow$ mIoU $\uparrow$
1 ICP-4D (Ours) 70.3 70.0 70.5
2 YAKD 68.6 70.8 66.5
3 hphnngcquan 68.5 70.4 66.7
4 KadirYilmaz 68.4 67.3 69.6
5 ConOcc 67.9 68.3 67.4
Appendix C Additional Experimental Results
Table A.3: Analysis on performance over the number of superimposed scans.
Memory (GB) Time (s) Validation set Test set
Method # Scan LSTQ $_{\text{assoc}} S$$_{\text{cls}}$ IoUSt IoUTh LSTQ $_{\text{assoc}} S$$_{\text{cls}}$ IoUSt IoUTh
(a) IoU-based Association
4D-PLS [2] 2 - - 59.9 58.8 61.0 65.0 63.1 - - - - -
4 8.0 3.48 62.7 65.1 60.5 65.4 61.3 56.9 56.4 57.4 66.9 51.6
black!10black 4D-Stop [20] 2 15.9 3.67 66.4 71.8 61.4 64.9 64.1 62.9 67.3 58.8 68.3 53.3
4 18.9 4.32 67.0 74.4 60.3 65.3 60.9 63.9 69.5 58.8 67.7 53.8
black!10black Eq-4D-Stop [63] 2 11.3 2.89 68.9 74.8 63.5 65.7 68.4 65.4 69.7 61.5 68.7 59.1
4 12.1 3.18 70.1 77.6 63.4 66.4 67.1 67.0 72.0 62.4 69.1 60.9
black!10black Mask4Former [54] 2 10.4 3.51 70.5 74.3 66.9 67.1 66.6 68.4 67.3 69.6 72.7 65.3
4 17.1 8.56 71.9 76.3 67.8 66.9 69.0 68.0 69.7 66.3 70.7 60.3
black!10black      (c) Detect & Track Association
ICP-4D (Ours) 1 3.2 2.35 72.7 78.3 67.5 65.2 70.6 70.3 70.0 70.5 72.7 67.4
C.1 Leaderboard Results
Table A.2 shows the SemanticKITTI [4] challenge leaderboard ranks on the hidden test set, ordered by LSTQ. In this leaderboard, we observe that our proposed ICP-4D outperforms not only all published methods but also every competition participant. Specifically, we achieve 1st place in LSTQ as of submission date. These results demonstrate the strong generalization capability and practical effectiveness of our method under challenging scenarios.
C.2 Additional Results
Table A.4: Additional comparisons across multiple metrics. ( ) denotes the number of input scans (i.e., 1f = single, 2f = two).
Method Metric MOTSA sMOTSA PTQ sPTQ
4D-StOP (2f) [20] 31.0 27.1 53.4 53.7
EQ-4D-StOP (2f) [63] 43.6 39.6 56.4 56.7
Mask4D (1f) [30] 52.1 48.1 58.4 58.5
Mask4Former (2f) [54] 47.3 42.1 48.3 48.5
black!10black Mask4Former (1f) - - - -
[…]
Table A.5: Comparison of the quantitative results on panoptic nuScenes [13].
Thing classes (AQ / IoU) Stuff classes (IoU only)
Model LSTQ Sassoc Scls
$\blacksquare$ barrier
$\blacksquare$ bicycle
$\blacksquare$ bus
$\blacksquare$ car
$\blacksquare$ const. veh.
$\blacksquare$ motorcycle
$\blacksquare$ pedestrian
$\blacksquare$ traffic cone
$\blacksquare$ trailer
$\blacksquare$ truck
$\blacksquare$ drive. surf.
$\blacksquare$ other flat
$\blacksquare$ sidewalk
$\blacksquare$ terrain
$\blacksquare$ manmade
$\blacksquare$ vegetation
Mask4Former [54] 67.6 63.0 72.5 48.3/76.9 42.6/37.0 60.5/90.2 76.7/92.7 42.7/44.9 61.1/77.0 53.5/74.0 48.9/55.7 38.7/47.8 62.2/77.8 95.6 69.2 71.5 74.6 88.5 87.0
Mask4D [30] 58.5 54.1 63.1 29.7/64.3 27.8/20.5 60.5/79.2 72.5/83.5 29.5/21.8 52.5/72.0 42.3/65.9 39.5/43.0 19.2/35.2 58.9/71.8 94.2 53.9 64.9 71.3 85.2 84.5
CA-Net [29] 63.4 56.6 71.0 30.1/74.5 31.9/29.1 56.5/88.5 73.3/91.2 38.6/43.2 63.2/78.7 53.6/71.2 34.6/54.6 39.5/48.3 62.1/77.3 95.6 68.4 70.3 72.7 86.7 85.2
ICP-4D (Ours) 67.9 65.0 71.0 40.7/74.5 44.1/29.1 58.9/88.5 80.9/91.2 40.1/43.2 65.9/78.7 67.2/71.2 52.0/54.6 40.6/48.3 65.7/77.3 95.6 68.4 70.3 72.7 86.7 85.2
Results on Multiple Metrics. We further present additional results with multiple metrics to enable a more comprehensive comparison. MOTSA [45] is widely used for evaluating multi-object tracking and segmentation. Although MOTSA emphasizes the recognition quality rather than temporal consistency, it nonetheless offers a straightforward indicator of the ID switches. Moreover, PTQ [17] is designed to jointly evaluate panoptic segmentation and tracking quality in a holistic manner. In Table C.2, evaluations are carried out on the SemanticKITTI [4] validation set. We observe that ICP-4D consistently surpasses all baselines across multiple metrics, including LSTQ. This indicates that the strength of ICP-4D goes beyond any particular metric, offering a robust and practically reliable solution for instance association.
C.3 Hyperparameter Analysis
In this section, we examine the impact of different matching thresholds on instance association performance. Figure A.1 illustrates the LSTQ scores of ICP-4D on the SemanticKITTI [4] validation set. We empirically determine the matching thresholds $\tau_{iou}$, $\tau_{dist}$, $\tau_{center}$, $\tau_{cov}$} over {0.1, 0.2, 0.3, 0.4, 0.5}. Here, we observe three insights. First, extremely low value of threshold $\tau_{iou}$ often yields decent performance due to over-association, whereas high thresholds lead to over-filtering and degraded matching quality. Second, excessively large $\tau_{center}$ and $\tau_{cov}$ cause static instances to be matched too liberally, leading to unreliable association. Lastly, our proposed ICP-4D delivers performance gains over all baselines, regardless of the hyperparameter choices.
Figure A.1: Hyperparameter Analysis.
C.4 Quantitative Results
We additionally provide the results of each semantic class on panoptic nuScenes [13] to facilitate a more comprehensive understanding. We present our reproduced results obtained from officially given training configurations. Note that 4D-Stop [20] and EQ-4D-Stop [63] do not provide official class-wise results for the panoptic nuScenes benchmark, whereas only overall metrics such as LSTQ are available. For this reason, class-wise comparisons against these methods are excluded to avoid confusion and ensure fairness. In Table A.5, our method yields overall performance improvements, particularly in safety-critical thing classes (e.g., bicycle, car, and pedestrian). Furthermore, classes with relatively fewer points, such as pedestrians and traffic cones, exhibit stronger association performance against the baselines. This observation aligns with our earlier analysis on SemanticKITTI [4] presented in the main manuscript.
C.5 Qualitative Results
In this section, we present additional qualitative analysis by illustrating the results of SemanticKITTI [4] and panoptic nuScenes [13] dataset. As clearly shown in Figure A.2 and  A.3, ICP-4D maintains robust ID consistency even in challenging scenarios. Our method explicitly incorporates the inherent geometric relations between instances and adopts a delicately designed pipeline that is resilient to occlusion and noisy predictions. In contrast, Mask4Former [54] associates instances by selecting the majority ID among spatially overlapped points, making it vulnerable to imprecise predictions and leading to frequent ID switches (see the 1st, 2nd, and 3rd rows of Figure A.2). Apart from this, Mask4D [30] struggles with separating nearby objects and accumulates errors from early mispredictions (see the 1st, 2nd, and 4th rows of Figure A.2). Furthermore, ICP-4D also operates well on the panoptic nuScenes dataset, which is sparse and large temporal gaps between consecutive scans. Specifically, in the 2nd and 4th rows of Figure A.3, we observe that the instances have no overlap and sparse representations. The baselines confuse the instance IDs due to their partial observations and rapid motions. However, ICP-4D successfully finds the corresponding instances even when they are distant. Also, from the 1st row of Figure A.3, we observe that ICP-4D can effectively associate instances in complex scene where pedestrians are heavily entangled.
Appendix D Discussion
D.1 Limitation and Future work
We identify several limitations of our proposed approach, which inherits issues from both pre-trained 3D panoptic models and LiDAR sensors. While ICP-4D demonstrates strong association performance with a cost efficient pipeline, real-time deployment remains challenging. Our method highly depends on the performance of the underlying 3D panoptic model; thus, its runtime can be further improved as point cloud processing continues to advance. In addition, partial observation of instances in LiDAR sensing make it difficult to reliably associate instances. Incorporating trajectory cues represents a promising direction to further stabilize instance association. Lastly, our pipeline forces the requirement of camera parameters as a prerequisite for transforming the point coordinates into the world coordinates. Relaxing this requirement would improve the applicability of our method in broader real-world settings. We hope these insights encourage continued progress toward scalable 4D perception systems.
Figure A.2: Qualitative comparison on semanticKITTI [4] validation set. We visualize the association results over five consecutive scans for both the baselines and our ICP-4D. Different colors represent different instances. The dotted boxes zoom in for clear comparison. For clarity, we use circled numbers (①–⑤) to represent the frame indices.
Figure A.3: Qualitative comparison on panoptic nuScenes [13]. We visualize the association results over five consecutive scans for both the baselines and our ICP-4D. Different colors represent different instances. The dotted boxes zoom in for clear comparison. For clarity, we use circled numbers (①–⑤) to represent the frame indices.